The effectiveness of a blood-pressure drug is being investigated. An experimenter finds that, on average, the reduction in systolic blood pressure is 37.4 for a sample of size 23 and standard deviation 11.2. Estimate how much the drug will lower a typical patient's systolic blood pressure (using a 90% confidence level). Assume the data is from a normally distributed population.

Enter your answer as a tri-linear inequality accurate to three decimal places.
Answer #1
sample mean 'x̄= 37.400
sample size   n= 23.00
sample std deviation s= 11.20
std error 'sx=s/√n= 2.3354
for 90% CI; and 22 df, value of t= 1.717
margin of error E=t*std error    = 4.010
lower bound=sample mean-E = 33.390
Upper bound=sample mean+E = 41.410
from above 90% confidence interval for population mean =(33.390 <μ < 41.410)